Operations Report December 13th

Crew 200 Operations Report 13-Dec-2018

SOL: 11

Name of person filing report: Antoine Bocquier

Non-nominal systems: generator autostart
Notes on non-nominal systems: None

Generator (hours run): From 5:30 pm last night to around 9 am today (SOC was still at 68% this morning). On tonight at 4:50pm.

Solar— SOC 77%
Diesel Reading – 100 % (refilled)
Propane Reading – 27 %
Ethanol Free Gasoline – N/A gallons

Water (auxillary tank) – 0 gallons
Water (static tank) – 340 gallons
Auxillary to Static tank transfer – no
Gallons transferred: 0
Water in GreenHab – high level
Water (loft) – Static to Loft Pump used – yes
Water Meter: 3

Toilet tank emptied: yes

Deimos rover used: no, still not functional
Hours: 141.5
Beginning charge:
Ending charge: 70%
Currently charging: no

Sojourner rover used: yes (by media)
Hours: 43
Beginning charge:
Ending charge: 100%
Currently charging: yes

Spirit rover used: yes (by media)
Hours: 62.1
Beginning charge:
Ending charge: 100%
Currently charging: yes

Opportunity rover used: yes (by media and us)
Hours: 42.9
Beginning charge: 100%
Ending charge: 72%
Currently charging: yes

Curiosity rover used: yes (by media)
Hours: 58.8
Beginning charge:
Ending charge: 100%
Currently charging: yes

Notes on rovers: Rovers were used today by the media, we used Opportunity for our planned EVA, whose batteries did pretty well.
# Hours the ATVs were Used today: 0
Notes on ATVs: ATVs were taken out of active service and are now parked away from the station

HabCar used and why, where? No
CrewCar used and why, where? No

General notes and comments: We repaired a tarp this morning, on the RAM tunnel. No issues to report with the heater.

Summary of internet: Nothing to report.

Summary of suits and radios: we performed an EVA this afternoon (we had to delay it as the media were recording this morning, approved by Shannon. Spacesuit 3 still not functional. On the one piece suits, it may be useful to have a chine guard to prevent the chine to hurt the bottom of the helmet especially when driving (mostly for short people).

Summary of Hab operations: repaired a tarp.

Summary of GreenHab operations: Regular activities

Summary of ScienceDome operations: Nothing to report.

Summary of RAM operations: Nothing to report, will be cleaned up.

Summary of any observatory issues: Musk Solar observatory – Nominal. Robotic Observatory – Nominal

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report